# Information Limits of Joint Community Detection and Finite Group Synchronization

### v1

@@ 4 Discussions
Algorithm design
From an algorithm perspective, since directly solving the MLE (1) is non-convex and computationally intractable, finding an efficient algorithm with a competitive performance becomes necessary in practice. In the context of community detection, efficient approaches based on semidefinite programming (SDP) [1, 37, 38, 4, 57, 47] and spectral method [2, 67, 46] have been proposed. In particular, they are known to achieve the information-theoretic limit (13) for exact recovery on the SBM [1, 2, 37, 57]. In light of this, similar algorithms [24, 23, 12, 66] have been applied to the joint problem, and it is natural to expect that these methods also work well all the way down to the information limit derived in this paper.
@@
Unfortunately, our result suggests that none of these methods are likely to reach the information limit, indicating a performance gap exists compared to the MLE (1). For instance, the SDP developed in [24] for solving the joint problem exhibits a sharp threshold for exact recovering both communities and group elements, which is given as
$a-\sqrt{2b}\log\left(\frac{ea}{\sqrt{2b}}\right)=2.$ (18)
as shown in Figure 3. Notably, the statistical model studied in [24] is slightly different to ours, where [24] focuses on $\mathcal{G}_{M}=\text{SO}(2)$ that is infinite, as apposed to the finite group considered in this work. Therefore, approximately, one may compare the threshold (18) by the SDP to the information limit (17) with a large $M$. As a result, the performance gap is clearly demonstrated.
To further understand the performance gap, our conjecture is that all existing approaches essentially differ from the MLE (1) and instead aim to solve an alternative optimization problem that allows (convex) relaxations to be smoothly applied. As an example, the authors in [24, 23] consider the orthogonal group $\textrm{O}(d)$ and design the following program for recovery:
@@
Giant component of Erdős–Rényi graphs
An important technical ingredient of our theory is analyzing the giant component of Erdős–Rényi graphs [20]. As aforementioned, it is well-known that an Erdős–Rényi graph $G(n,p)$ of size $n$ and connection probability $p=\frac{a\log n}{n}$ for some constant $a$ is connected with high probability if and only if $a>1$. Furthermore, when $a<1$ such as the graph is disconnected, a giant component of size close to $n$ is expected to exist with high probability. To be specific, let $Z_{n}$ denote the size of the largest connected component in the graph, a common result [31, Theorem 2.14] is
$\mathcal{P}\left(Z_{n}>\left(1-n^{-a+\epsilon+o(1)}\right)n\right)>1-n^{-\epsilon}$ (20)
which holds for any $\epsilon<a$. However, (20) is too loose to be applied for our analysis in Section 5.2. Therefore, we improve (20) with the following sharp result obtained
@@
which is necessary for acquiring the sharp threshold (12), see Theorem A.4 in Appendix A for the details. Remarkably, (21) cannot be derived by the conventional moment methods such as Markov’s or Chebyshev’s inequality, which are generally applied in random graph theory (e.g. see [7]). Instead, we prove (21) by conducting a delicate combinatorial analysis, which serves as one of the main technical challenges in our proof.
Future works
Our analysis presented in this work can be generalized in several aspects, listed as follows:
@@
    Although we restrict our attention on the simplest scenario of two equal-sized communities, our proof framework can be naturally extended to more generalized cases of multiple clusters with different sizes
@@
    Depending on the applications, one may consider a more involved statistical model such as the one studied in [23] where the group transformation $g_{ij}$ observed on an edge $(i,j)$ within the same community is perturbed by noise rather than being exact i.e., $g_{ij}=g_{i}^{*}\left(g_{j}^{*}\right)^{-1}$.
@@
    Based on our discussion in Section 4, finding an efficient algorithm that achieves the information limit is still an open problem. Here, our result suggests that the key might be focusing on the recovery of communities which is the prerequisite for recovering group elements.
5 Proof of the theorems
In this section, we describe the proofs of our theorems presented in Section 3. For ease of presentation, some technical details are deferred to the Appendices.
5.1 Preliminaries
@@
Four subsets.
Under the setting of two equal-sized communities, given the ground truth of the community assignment $\bm{\kappa}^{*}$ and any alternative hypothesis $\bm{\kappa}$, we define the following four subsets of nodes, which will be frequently used in our proof:
@@
For instance, $S_{12}$ stands for the set of nodes that is assigned to $S_{1}$ in $\bm{\kappa}$ but actually belongs to $S_{2}^{*}$ in $\bm{\kappa}^{*}$. See Figure 4 for an illustration. Notably, all the four subsets are mutually disjoint and their union covers all nodes in the network. Also, one can immediately observe the following equations:
@@
which hold for any valid hypothesis $\bm{\kappa}$.
Derivation of the MLE (1)
We start by expanding the likelihood $\mathcal{P}(\mathcal{E},\mathfrak{G}|\bm{\kappa},\bm{g})$ as
@@
$=\mathcal{P}(\mathcal{E}|\bm{\kappa})\mathcal{P}(\mathfrak{G}|\bm{\kappa},\bm{g},\mathcal{E})$
which is separated into the likelihood of edge connections $\mathcal{P}(\mathcal{E}|\bm{\kappa})$ and the one of group transformations given the edges, i.e. $\mathcal{P}(\mathfrak{G}|\bm{\kappa},\bm{g},\mathcal{E})$. According to the statistical model, $\mathcal{P}(\mathcal{E}|\bm{\kappa})$ is given as
@@
$\propto\left(\frac{p(1-q)}{q(1-p)}\right)^{|\mathcal{E}_{\text{inner}}|}$
where ${\mathcal{E}_{\text{inter}}}$ denotes the set of edges across different communities, as opposed to $\mathcal{E}_{\text{inner}}$ defined in (2), and we use the fact that $|\mathcal{E}|=|\mathcal{E}_{\text{inner}}|+|\mathcal{E}_{\text{inter}}|$. Also, for $\mathcal{P}(\mathfrak{G}|\bm{g},\bm{\kappa},\mathcal{E})$, we have
@@
$\propto\mathbbm{1}\left(g_{ij}=g_{i}g_{j}^{-1};\;\forall(i,j)\in\mathcal{E}_{\text{inner}}\right)\cdot M^{|\mathcal{E}_{\text{inner}}|}$
@@
Given the above, by plugging (25) and (26) into (24), the MLE can be written as
@@
s.t. $\quad n_{k}=n_{k}^{*},\quad k=1,\ldots,K$
where we apply the assumption that all cluster sizes are given.
At first glance, solving (27) requires the prior knowledge of the model parameters $p,q$, and $M$. However, by assuming $Mp(1-q)>q(1-p)$ or $Mp(1-q)\leq q(1-p)$, and making the indicator function as constraints, (27) becomes (1) which does not involve any model parameter specifically.
Likelihood ratio
Given the true parameter $\bm{x}^{*}=\{\bm{\kappa}^{*},\bm{g}^{*}\}$ and a hypothesis $\bm{x}=\{\bm{\kappa},\bm{g}\}$, our proof is based on the likelihood ratio defined as
@@
To this end, by using the result in (25), we get
$\frac{\mathcal{P}(\mathcal{E}|\bm{\kappa})}{\mathcal{P}(\mathcal{E}|\bm{\kappa}^{*})}=\left(\frac{Mp(1-q)}{q(1-p)}\right)^{|\mathcal{E}_{\text{inner}}|-|\mathcal{E}_{\text{inner}}^{*}|}$ (28)
where $\mathcal{E}_{\text{inner}}^{*}$ denotes the set of in-cluster edges according to $\bm{\kappa}^{*}$. Furthermore, recall the four subsets defined in (22), then notice that
$|\mathcal{E}_{\text{inner}}|-|\mathcal{E}_{\text{inner}}^{*}|$ $=|\mathcal{E}(S_{11},S_{12})|+|\mathcal{E}(S_{21},S_{22})|-|\mathcal{E}(S_{11},S_{21})|-|\mathcal{E}(S_{12},S_{22})|$ (29)
$=r-r^{*}$
@@
One can interpret $r$ (resp. $r^{*}$) as the number of connections that are within the same community in $\bm{\kappa}$ (resp. $\bm{\kappa}^{*}$) but across different ones in $\bm{\kappa}^{*}$ (resp. $\bm{\kappa}$). See Figure 4 for an illustration. Also, by using (26), we obtain
$\frac{\mathcal{P}(\mathfrak{G}|\bm{\kappa},\bm{g},\mathcal{E})}{\mathcal{P}(\mathfrak{G}|\bm{\kappa}^{*},\bm{g}^{*},\mathcal{E})}=\frac{\mathbbm{1}\left(g_{ij}=g_{i}g_{j}^{-1};\;\forall(i,j)\in\mathcal{E}_{\text{inner}}\right)}{\mathbbm{1}\left(g_{ij}=g_{i}^{*}(g_{j}^{*})^{-1};\;\forall(i,j)\in\mathcal{E}_{\text{inner}}^{*}\right)}.$
Combining the results above yields the likelihood ratio
$\frac{\mathcal{P}(\bm{y}|\bm{x})}{\mathcal{P}(\bm{y}|\bm{x}^{*})}$ $=\left(\frac{Mp(1-q)}{q(1-p)}\right)^{r-r^{*}}\cdot\frac{\mathbbm{1}\left(g_{ij}=g_{i}g_{j}^{-1};\;\forall(i,j)\in\mathcal{E}_{\text{inner}}\right)}{\mathbbm{1}(g_{ij}=g_{i}^{*}(g_{j}^{*})^{-1};\;\forall(i,j)\in\mathcal{E}_{\text{inner}}^{*})}$ $=r-r^{*}$ (31)
@@
$g_{i}^{*}=g_{0},\quad\forall i\in\{1,\ldots,n\}$ (32)
where $g_{0}$ stands for the neutral (zero) element of the group $\mathcal{G}_{M}$. Also, we denote
$\mathcal{P}_{\bm{x}^{*}}(\cdot):=\mathcal{P}(\cdot|\bm{x}^{*}),\quad\mathcal{P}_{\bm{x}}(\cdot):=\mathcal{P}(\cdot|\bm{x})$
@@
Step 1: A union bound

### v2

@@ 4 Discussions
4.1 Existing approaches and their limits
From an algorithm perspective, since directly solving the MLE (1) is non-convex and computationally intractable, finding an efficient algorithm with a competitive performance becomes necessary in practice. In the context of community detection, efficient approaches based on semidefinite programming (SDP) [2, 38, 39, 5, 60, 50] and spectral method [3, 72, 48] have been proposed. In particular, they are known to achieve the information-theoretic limit (11) for exact recovery on the SBM [2, 3, 38, 60]. In light of this, similar algorithms [25, 24, 13, 70] have been applied to the joint problem, and it is natural to expect that these methods also work well all the way down to the information-theoretic limit derived in this paper.
@@ 4.1 Existing approaches and their limits
Unfortunately, our result suggests that none of these methods can reach the information-theoretic limit, indicating that there is a performance gap compared to MLE (1). For instance, the SDP developed in [25] for solving the joint problem exhibits a sharp threshold for the exact recovery of both communities and group elements, which is given as
$a-\sqrt{2b}\log\left(\frac{ea}{\sqrt{2b}}\right)=2,$ (22)
as shown in Figure 3. Notably, the statistical model studied in [25] is slightly different to ours, where [25] focuses on the compact Lie group $\text{SO}(2)$, as opposed to the finite group considered in this work. Therefore, approximately, one may compare the threshold (22) by the SDP to the information-theoretic limit (21) with a large $M$. As a result, the performance gap is clearly demonstrated by comparing Figure 3 with Figure 2.
For spectral method-based approaches [24, 13, 70, 7], it is not surprising to observe similar performance gaps exist as well. For example, [13] proposes using generalized power method and provides the corresponding condition for exact recovery [13, Theorem 2]. That is, under the setting of compact Lie group $\text{SO}(2)$ and two equal-sized clusters, the sufficient condition for exact recovery is:
$\sqrt{2b}<a\quad\text{and}\quad a-\sqrt{2b}\log\left(\frac{ea}{\sqrt{2b}}\right)>2,$ (23)
which is almost identical to (22) by SDP with the extra requirement $\sqrt{2b}<a$, indicating a performance gap similar to the SDP [25] exists. In a same spirit of methodology, [24] employs QR-decomposition for recovery and develops the condition for exact recovery [24, Theorem 5.2] under the same setting (two equal-sized clusters and compact Lie group $\text{SO}(2)$) as222 The original condition provided in [24, Theorem 5.2] is written as $\frac{\sqrt{\left(p(1-p)+q\right)\log(2n)}}{p\sqrt{n}}\leq c_{0}$ and (24) is obtained by plugging the scaling $p=\frac{a\log n}{n}$ and $q=\frac{b\log n}{n}$.
$\frac{\sqrt{a+b}}{a}\leq c_{0},$ (24)
for some unknown constant $c_{0}>0$. While the performance gap may not be explicitly evident from (24) due to the unknown value of $c_{0}$, it is clearly illustrated in the experimental results shown in [24, Figure 2]. Besides, the performance of another spectral method-based solution [7] is examined in [24, Figure 5], revealing a phase transition threshold similar to [24]. Lastly, an improved approach based on [24] by aggregating multiple irreducible representations of group transformations is presented in [70]. Despite this refinement, the method exhibits a similar performance trend (see e.g. [70, Figure 3]) as [24], and thus, the performance gap for algorithms with polynomial time complexity persists.
4.2 Understanding the performance gap
To further understand the performance gap, note that all the aforementioned convex relaxation-based approaches exhibits a common pattern: recovery becomes more challenging as $q$ (the probability of connection across communities) or $b$ increases. This behavior contrasts with the information-theoretic limits displayed in Figure 2, where recovery can actually become easier as $q$ increases. To interpret this, as discussed in Section 3.1, inter-cluster edges in community detection typically carry no useful information but act purely as noise. However, in the joint problem setting, where additional group transformations are observed, inter-cluster edges can aid recovery by helping to eliminate incorrect hypothesis that violate group transformation consistency. Therefore, the performance gap arises from the inability of these methods to fully exploit the potential advantages offered by such “noisy” edges.
In addition, from the perspective of problem formulation, it is evident that all existing approaches deviate from the original MLE (1). Instead, they are designed to solve alternative optimization problems that are more amenable to (convex) relaxation techniques. As an example, the authors in [25, 24] consider the orthogonal group $\textrm{O}(d)$ and design the following program for recovery:
@@ 4.2 Understanding the performance gap
To bridge the performance gap between existing algorithms and the information-theoretic limits, our earlier discussion suggests that a more refined approach could be developed by carefully verifying cycle consistency – similar to our proof technique used in Section 6.3 for Theorem 3.1. However, such an algorithm would likely entail significantly higher computational cost compared to existing convex relaxation-based methods. For instance, similar idea of checking cycle consistency has been recently applied to group synchronization in [49], where the proposed algorithm is of complexity of $O(Ln^{L})$ (see [49, Section 4.2.7]), with $L$ denoting the cycle length being checked. To manage the computational cost, it is desirable to use the shortest cycle in the graph (e.g. $L=3$). However, this becomes problematic in sparse graphs (which is often the case in practice) where nodes typically participate only in long cycles with $L\gg 3$. For example, in an Erdős–Rényi graph $G(n,p)$ with $p=\Theta(\frac{\log n}{n})$, one can show that for any node, the expected number of cycles with length $L$ passing through it is [^4] Here, (${n-1\choose L-1}$ determines the set of nodes involved in the cycle, $\frac{(L-1)!}{2}$ stands for the order of nodes traversing along the cycle, and $p^{L}$ ensures the cycle are indeed connected.
(${n-1\choose L-1}\cdot\frac{(L-1)!}{2}\cdot p^{L}=O\left(\frac{(\log n)^{L}}{n}\right).$
Therefore, a typical cycle length would be $L=\omega\big{(}\frac{\log n}{\log\log n}\big{)}$, indicating the scarcity of short cycles with $L=\Theta(1)$ in the sparse regime. As a result, the necessity of checking long cycles makes the algorithm inefficient and less favorable comparing to the existing convex relaxation based approaches.
Moreover, the algorithm that checks cycle-consistency relies on the assumption that the measurements $g_{ij}$ are perfect for intra-cluster connections, whereas in practice these are typically noisy. As a result, this fact further limits the practicality of cycle-consistency checks. In contrast, all the existing methods tolerate moderate noise as long as it stays below a certain threshold. Given the above, designing an efficient and robust algorithm that can handle both sparse structures and noisy measurements remains an open and challenging problem.
4.3 Giant component of Erdős–Rényi graphs
An important technical ingredient of our theory is analyzing the giant component of Erdős–Rényi graphs [21]. As aforementioned, it is well-known that an Erdős–Rényi graph $G(n,p)$ of size $n$ and connection probability $p=\frac{a\log n}{n}$ for some constant $a$ is connected with high probability if and only if $a>1$. Furthermore, when $a<1$ such as the graph is disconnected, a giant component of size close to $n$ is expected to exist with high probability. To be specific, let $Z_{n}$ denote the size of the largest connected component in the graph. By following standard techniques commonly used in the study of random graphs (e.g., [32, Theorem 2.14])555[32, Theorem 2.14] focuses on the regime where $p=\Theta(1/n)$, whereas our interest lies in the denser regime $p=\Theta(\log n/n)$. Therefore, we cannot directly apply [32, Theorem 2.14], but we adopt the same proof techniques to derive the bound in (26). In particular, we first show that there is no component of intermediate size (i.e., between $\beta_{1}\log n$ and $\beta_{2}n$ for some constants $\beta_{1},\beta_{2}>0$), and then we upper bound the total number of vertices in small components (those of size less than $\beta_{1}\log n$.), we can derive the following result regarding $Z_{n}$,
$\mathcal{P}\left(Z_{n}>\left(1-n^{-a+\epsilon+o(1)}\right)n\right)>1-n^{-\epsilon},$ (26)
which holds for any $\epsilon<a$. However, (26) is too loose to be directly applied in our analysis in Section 6. To address this, we improve upon (26) by establishing the following sharper result:
@@ 4.3 Giant component of Erdős–Rényi graphs
which is essential for deriving the sharp threshold in (10); see Theorem A.4 in Appendix A for details. Notably, the bound in (27) cannot be obtained via conventional moment methods such as Markov’s or Chebyshev’s inequalities, which are commonly used in random graph theory (e.g., see [8]). Instead, we establish (27) through a delicate combinatorial analysis, which constitutes one of the main technical challenges of our proof.
4.4 Future works
In summary, our analysis presented in this work can be generalized in several aspects, listed as follows:
@@ 4.4 Future works
    Although we restrict our attention on the simplest scenario of two equal-sized communities, our proof framework can be naturally extended to more generalized cases of multiple clusters with different sizes. In such scenarios, we expect that the same proof techniques – particularly our key technical ingredients such as the cycle consistency analysis in Section 6.3 – can still be applied, though with additional care to account for the presence of multiple communities. As $K$ increases, exact recovery is expected to become more challenging due to a higher number of inter-cluster edges.
@@
    Our theory assumes the group transformation $g_{ij}$ is perfectly observed on connections within the same community. However, it is more common in practice where $g_{ij}$ is perturbed by noise. Therefore, in order to extend our analysis to account for such noise, additional care must be taken in the cycle consistency analysis (e.g., Section 6.3 for the upper bound). Specifically, one should not expect exact cycle consistency, and the error in group alignment may accumulate in long cycles. The presence of inconsistency may no longer reliably indicate the existence of inter-cluster edges within the cycle.
@@
    Another promising direction for future work is to establish the theoretical guarantees for weak (partial) recovery, which is important in many practical scenarios where partial recovery may already suffice for downstream applications. However, the proof techniques required for weak recovery may differ significantly from those used in our current analysis for exact recovery. Some insights may be drawn from existing studies on weak recovery in community detection (e.g., [1, 18, 56]), which show that weak recovery is achievable in the sparse regime $p=a/n$, $q=b/n$ for some constants $a,b$, provided that giant components exist within communities.
  5. 5.
    Based on our discussion in Section 4.2, finding an efficient algorithm that achieves the information-theoretic limit remains an open problem. Our results suggest focusing on the recovery of communities, which is a prerequisite to accurately recover group elements. Moreover, developing an efficient strategy to check the cycle consistency could be the key to significantly improving the performance of existing approaches.
5 Proof Preliminaries
@@
5.1 Four subsets
Under the setting of two equal-sized communities, given the ground truth of the community assignment $\bm{\kappa}^{*}$ and any alternative hypothesis $\bm{\kappa}$, we define the following four subsets of nodes, which greatly facilitate our analysis and thus will be frequently used in our proof:
@@ 5.1 Four subsets
For instance, $S_{11}$ denotes the set of nodes that are correctly identified as belonging to cluster 1, while $S_{12}$ stands for the set of nodes that are incorrectly assigned to cluster 1 but actually belongs to cluster 2. See Figure 4 for an illustration. Notably, the four subsets are mutually disjoint and their union covers all nodes in the network. Also, one can immediately observe the following equations:
@@
which hold for any valid hypothesis $\bm{\kappa}$ under the assumption that the two communities are of equal size.
5.2 Derivation of the MLE (1)
Our theory is built based on the MLE program (1) where we start by expanding the likelihood $\mathcal{P}(\mathcal{E},\mathfrak{G}|\bm{\kappa},\bm{g})$ as
@@ 5.2 Derivation of the MLE (1)
$=\mathcal{P}(\mathcal{E}|\bm{\kappa})\mathcal{P}(\mathfrak{G}|\bm{\kappa},\bm{g},\mathcal{E}),$
which is the product of the likelihood of edge connections $\mathcal{P}(\mathcal{E}|\bm{\kappa})$ and the likelihood of group transformations given the edges, i.e., $\mathcal{P}(\mathfrak{G}|\bm{\kappa},\bm{g},\mathcal{E})$.
According to the statistical model, $\mathcal{P}(\mathcal{E}|\bm{\kappa})$ is given as
@@
$\propto\left(\frac{p(1-q)}{q(1-p)}\right)^{|\mathcal{E}_{\text{inner}}|},$
where ${\mathcal{E}_{\text{inter}}}:=\{(i,j)\;|\;(i,j)\in\mathcal{E},\;i<j,\;\kappa_{i}\neq\kappa_{j}\}$ denotes the set of edges across different communities, as opposed to $\mathcal{E}_{\text{inner}}$ defined in (2) for the set of edges within the same community, and we have used the fact that $|\mathcal{E}|=|\mathcal{E}_{\text{inner}}|+|\mathcal{E}_{\text{inter}}|$.
For $\mathcal{P}(\mathfrak{G}|\bm{g},\bm{\kappa},\mathcal{E})$, we have
@@
$\propto\mathbbm{1}\left(g_{ij}=g_{i}g_{j}^{-1};\;\forall(i,j)\in\mathcal{E}_{\text{inner}}\right)\cdot M^{|\mathcal{E}_{\text{inner}}|},$
@@
Given the above, by plugging (31) and (32) into (30), the MLE problem can be written as
@@
s.t. $\quad n_{k}=n_{k}^{*},\quad k=1,\ldots,K,$
where we apply the assumption that the size of each community is provided, and note that (33) applies to arbitrary community settings. At first glance, solving (33) requires the prior knowledge of the model parameters $p,q$, and $M$. However, by splitting (33) into the two scenarios $Mp(1-q)>q(1-p)$ or $Mp(1-q)\leq q(1-p)$ and making the indicator function as constraints, (33) becomes the aforementioned MLE forumlation (1), which does not involve any model parameter specifically.
5.3 Likelihood ratio
Given the true parameter $\bm{x}^{*}=\{\bm{\kappa}^{*},\bm{g}^{*}\}$ and a hypothesis $\bm{x}=\{\bm{\kappa},\bm{g}\}$, our proof is based on analyzing the likelihood ratio defined as
@@ 5.3 Likelihood ratio
Then, by using the likelihood derived in (31), we obtain
$\frac{\mathcal{P}(\mathcal{E}|\bm{\kappa})}{\mathcal{P}(\mathcal{E}|\bm{\kappa}^{*})}=\left(\frac{p(1-q)}{q(1-p)}\right)^{|\mathcal{E}_{\text{inner}}|-|\mathcal{E}_{\text{inner}}^{*}|},$ (34)
where $\mathcal{E}_{\text{inner}}^{*}$ denotes the set of in-cluster edges according to $\bm{\kappa}^{*}$, analogous to $\mathcal{E}_{\text{inner}}$ defined in (2) for the hypothesis $\bm{\kappa}$. Furthermore, under the setting of two equal-sized clusters and recall the four subsets defined in (28), we have
$|\mathcal{E}_{\text{inner}}|$ $=|\mathcal{E}(S_{11})|+|\mathcal{E}(S_{12})|+|\mathcal{E}(S_{21})|+|\mathcal{E}(S_{22})|+|\mathcal{E}(S_{11},S_{12})|+|\mathcal{E}(S_{21},S_{22})|,$
$|\mathcal{E}^{*}_{\text{inner}}|$ $=|\mathcal{E}(S_{11})|+|\mathcal{E}(S_{12})|+|\mathcal{E}(S_{21})|+|\mathcal{E}(S_{22})|+|\mathcal{E}(S_{11},S_{21})|+|\mathcal{E}(S_{12},S_{22})|.$
We suggest readers resorting to Figure 4 for a better understanding. Then, $|\mathcal{E}_{\text{inner}}|-|\mathcal{E}_{\text{inner}}^{*}|$ in (34) can be expressed as
$|\mathcal{E}_{\text{inner}}|-|\mathcal{E}_{\text{inner}}^{*}|$ $=\underbrace{|\mathcal{E}(S_{11},S_{12})|+|\mathcal{E}(S_{21},S_{22})|}_{=:r}-\underbrace{\left(|\mathcal{E}(S_{11},S_{21})|+|\mathcal{E}(S_{12},S_{22})|\right)}_{=:r^{*}}$ (35)
$=:r-r^{*},$
@@
One can interpret $r$ (resp. $r^{*}$) as the number of connections that are within the same community in $\bm{\kappa}$ (resp. $\bm{\kappa}^{*}$) but across different ones in $\bm{\kappa}^{*}$ (resp. $\bm{\kappa}$). See Figure 4 for an illustration.
Similarly, for the other part of the likelihood ratio, by using (32), we obtain
$\frac{\mathcal{P}(\mathfrak{G}|\bm{\kappa},\bm{g},\mathcal{E})}{\mathcal{P}(\mathfrak{G}|\bm{\kappa}^{*},\bm{g}^{*},\mathcal{E})}$ $=\frac{\mathbbm{1}\left(g_{ij}=g_{i}g_{j}^{-1};\;\forall(i,j)\in\mathcal{E}_{\text{inner}}\right)}{\mathbbm{1}\left(g_{ij}=g_{i}^{*}(g_{j}^{*})^{-1};\;\forall(i,j)\in\mathcal{E}_{\text{inner}}^{*}\right)}\cdot M^{\left(|\mathcal{E}_{\text{inner}}|-|\mathcal{E}_{\text{inner}}^{*}|\right)}$
$=\frac{\mathbbm{1}\left(g_{ij}=g_{i}g_{j}^{-1};\;\forall(i,j)\in\mathcal{E}_{\text{inner}}\right)}{\mathbbm{1}\left(g_{ij}=g_{i}^{*}(g_{j}^{*})^{-1};\;\forall(i,j)\in\mathcal{E}_{\text{inner}}^{*}\right)}\cdot M^{r-r^{*}},$
where we have applied (35) in the last step.
Now, by combining the results above, we can express the likelihood ratio as
$\frac{\mathcal{P}(\bm{y}|\bm{x})}{\mathcal{P}(\bm{y}|\bm{x}^{*})}$ $=\left(\frac{Mp(1-q)}{q(1-p)}\right)^{r-r^{*}}\cdot\frac{\mathbbm{1}\left(g_{ij}=g_{i}g_{j}^{-1};\;\forall(i,j)\in\mathcal{E}_{\text{inner}}\right)}{\mathbbm{1}(g_{ij}=g_{i}^{*}(g_{j}^{*})^{-1};\;\forall(i,j)\in\mathcal{E}_{\text{inner}}^{*})},$ (37)
@@
$g_{i}^{*}=I,\quad\forall i\in\{1,\ldots,n\},$ (38)
where $I$ stands for the identity element of group $\mathcal{G}_{M}$, i.e., $I\cdot g=g$. Also, we denote
$\mathcal{P}_{\bm{x}^{*}}(\cdot):=\mathcal{P}(\cdot|\bm{x}^{*}),\quad\mathcal{P}_{\bm{x}}(\cdot):=\mathcal{P}(\cdot|\bm{x}),$
@@
For readability, we split our proof into four major steps.
6.1 Step 1: A union bound